OpenStudy (anonymous):

combine into a single expression how do i do this 3/x+4/y

OpenStudy (anonymous):

Common denominators...just like if it were numbers. If you had 2/3 + 1/2 What would you do?

OpenStudy (agent0smith):

Get common denominators for\[\Large \frac{ 3 }{ x }+\frac{ 4 }{ y }\]multiply the left one by y/y, and the right one by x/x, to get a common denominator xy \[\Large \frac{ 3 }{ x }*\frac{ y }{ y }+\frac{ 4 }{ y }*\frac{ x }{ x }\]

OpenStudy (anonymous):

3y/xy+4x/yx? so it would be 3y+4x/xy?

jhonyy9 (jhonyy9):

yes just cross multiplie

jhonyy9 (jhonyy9):

Ok ?

jhonyy9 (jhonyy9):

so because the common denominator is xy ok ?

jhonyy9 (jhonyy9):

and will get (3y+4x)xy

OpenStudy (skullpatrol):

You need to write your answer like this: (3y+4x)/xy not 3y+4x/xy, or use $$\frac{3y+4x}{xy}$$

OpenStudy (anonymous):

i understand what all of you arre saying but i never totally understood the concept of common denominators

OpenStudy (anonymous):

actuallly i get it thanks alot everybody :D its means alot!!!

OpenStudy (agent0smith):

Just look for what is "missing" in each denominator, to get a common denominator.

OpenStudy (anonymous):

u mean like how in one we had x and the other we had y but thy werent joined?

OpenStudy (agent0smith):

Yes, on one denominator you only had an x, so it's missing a y. In the other, you only had a y, so it's missing an x, so to get them the same, you have to make the denominator xy for both.

OpenStudy (anonymous):

okay that makes sense thank you so much :)